OpenStudy (anonymous):

write the equation of the line which passes through (2,1) and is perpendicular to x=2? need the steps

OpenStudy (saifoo.khan):

y = 2

OpenStudy (anonymous):

how did you get that ?

OpenStudy (saifoo.khan):

look, @ x = 2 there is a line parallel to y-axis. whn u r going to draw a perp. line @ y = 2, the slope will be zero. 2 = 1(0) + c c = 2 so, the line is going to cut the y axis @ 2. so they equation will be: y =2

OpenStudy (anonymous):

hmm, isnt it y=1 :P

OpenStudy (anonymous):

2 = 1(0) + c assuming this is meant to equate it with y=mx+b for you got x and y mixed :|

OpenStudy (saifoo.khan):

but 1 x 0 = 0

OpenStudy (anonymous):

:|